### v1

# C:/Users/LENOVO/Dropbox/tex/bibfile.bib

Abstract.
Komjáth, Milner, and Polat investigated when a finitary matroid admits a partition into circuits. They defined the class of “finite matching extendable” matroids and showed in their compactness theorem that those matroids always admit such a partition. Their proof is based on Shelah’s singular compactness technique and a careful analysis of certain $\triangle$-systems.
We provide a short, simple proof of their theorem. Then we show that a finitary binary oriented matroid can be partitioned into directed circuits if and only if, in every cocircuit, the cardinality of the negative and positive edges are the same. This generalizes a former conjecture of Thomassen, settled affirmatively by the second author, about partitioning the edges of an infinite directed graph into directed cycles. As side results, a Laviolette theorem for finitary matroids and a Farkas lemma for finitary binary oriented matroids are proven. An example is given to show that, in contrast to finite oriented matroids, ‘binary’ cannot be omitted in the latter result.
2020 Mathematics Subject Classification:
Primary: 05B35 Secondary: 03E05, 05A18, 05C63
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Grant No. 513023562 and partially by NKFIH OTKA-129211
1. Introduction
Decomposing complex structures into simple parts is a common theme in mathematical research, extending beyond combinatorics. Cycles are one of the most fundamental graph classes, leading to a natural question: When is it possible to partition a graph into cycles? For edge-partitions of finite graphs, the answer is straightforward. On the one hand, each cycle has an even contribution to the degree of each vertex, so the exclusion of vertices with odd degrees is necessary. On the other hand, if all degrees are even, a desired edge-partition can be constructed “greedily”. The problem becomes significantly harder for infinite graphs, especially uncountable ones. Having no vertices of odd degree is still necessary but insufficient. For example, in a two-way infinite path, each vertex has degree two, but it does not even contain a single cycle. Nash-Williams proved in his seminal paper [nash1960decomposition, p. 235 Theorem 3] that excluding odd cuts [^1] Finite cuts with an odd number of edges. is sufficient (and obviously necessary). New, simpler proofs of his theorem were found by L. Soukup [soukup2011elementary, Theorem 5.1] and Thomassen [thomassen2017nash]. In the same paper, Thomassen conjectured that the edges of a directed graph can be partitioned into directed cycles if and only if, in every cut, the cardinality of the edges going in one direction equals the cardinality of the edges going in the opposite direction. This was confirmed by the second author [joo2021dicycleDecomp, Theorem 2].
Welsh observed [welsh1969euler] that a finite binary matroid [^2] Definitions and basic facts about finitary matroids are given in Subsection 2.2. can be partitioned into circuits if and only if it has no odd cocircuit, which generalizes the analogous observation about graphs. Polat extended this to finitary binary matroids [polat1987compactness, Theorem 2.1 (iv) and (vi)], providing a common generalization of Nash-Williams’ theorem and Welsh’s result. Later, together with Komjáth and Milner [komjath1988compactness, Theorem 1], he further generalized his result by proving that matroids they call “finite matching extendable” (f.m.e.) can always be partitioned into circuits. Since binary matroids that not containing odd cocircuits are f.m.e., this is indeed a generalization. Their proof is based on Shelah’s singular compactness technique and a careful analysis of certain $\triangle$-systems. We begin this paper by providing a short and relatively simple proof of their theorem (Theorem 3.1).
A natural question is whether it is also possible to generalize the directed cycle edge-partition result [joo2021dicycleDecomp, Theorem 2] to finitary binary oriented matroids, as Polat generalized Nash-Williams’ theorem to finitary binary matroids. The first difficulty corresponds to the so-called Farkas lemma. In a directed graph, it is easy to see that every directed edge is contained in a directed cycle or a directed cut. This is known to be true for finite oriented matroids, but we will show that it may fail badly in infinite ones (see Claim 5.1). However, we prove that it holds true in the binary case (see Lemma 5.2). Using this, we prove that a finitary binary oriented matroid can be partitioned into directed circuits if and only if, in every cocircuit, the cardinality of the negative and positive edges are the same (see Theorem 6.1).
Let $\lambda$ be an infinite cardinal. An edge-partition ${G_{i}:i\in I}$ of a graph $G$ into subgraphs of size at most $\lambda$ is called $\lambda$-bond faithful if in each $G_{i}$ every bond of size less than $\lambda$ is also a bond of $G$, and each bond of $G$ of size at most $\lambda$ is a bond of a suitable $G_{i}$. Laviolette proved that an $\aleph_{0}$-bond faithful partition always exists, and, under the Generalized Continuum Hypothesis, $\lambda$-bond faithful partitions exist for every $\lambda$ (see [laviolette2005decompositions, Proposition 3]). L. Soukup showed that the Generalized Continuum Hypothesis can be omitted. We obtain a generalization of this result to finitary matroids (see Theorem 4.1).
The paper is structured as follows. In the next section, we survey the basic facts and notation related to set theory and matroid theory that we will use. Sections 3 and 4, due to the second author, discuss the new short proof of the compactness theorem by Komjáth et al. and the generalization of Laviolette’s theorem. Sections 5 and 6 are based on previous works of the first author with Carmesin, who kindly gave us permission to use them. Their respective contents are the generalization of the Farkas lemma and the directed circuit partition result. The framework of our proofs involves cutting up uncountable matroids by a chain of elementary submodels, a method popularized by L. Soukup in [soukup2011elementary].
2. Notation and preliminaries
2.1. Set theory
We use variables $\alpha$ and $\beta$ to stand for ordinal numbers, while $\kappa$, $\lambda$ and $\Theta$ denote cardinals. We write $\bigcup X$ for the union of the sets in $X$. A function $f$ is a set of ordered pairs such that if $\left\langle x,y\right\rangle,\left\langle x,z\right\rangle\in f$, then $y=z$. If $X$ is any set, (not necessarily $X\subseteq\mathsf{dom}(f)$), then we write $f\!\!\upharpoonright\!\!X$ for $\{\left\langle x,y\right\rangle\in f:\ x\in X\}$. The transitive closure of a set $X$ is the set whose elements are: the elements of $X$, the elements of the elements of $X$ etc. For a cardinal $\varTheta$, $H(\varTheta)$ denotes the set of those sets $X$ whose transitive closure is smaller than $\Theta$. We write $\mathcal{E}\prec H(\varTheta)$ if $(\mathcal{E},\in)$ is an elementary submodel of the first order structure $(H(\Theta),\in)$, i.e. $\emptyset\neq\mathcal{E}\subseteq H(\varTheta)$ and for every formula $\varphi(v_{0},\dots,v_{n-1})$ in the language of set theory and for every $x_{0},\dots,x_{n-1}\in\mathcal{E}$: $\varphi(x_{0},\dots,x_{n-1})$ holds true in the structure $(\mathcal{E},\in)$ iff it holds true in $(H(\Theta),\in)$. For a gentle introduction to elementary submodels and their combinatorial applications we refer to [soukup2011elementary].
Fact 2.1 ([soukup2011elementary, Claim 3.2]).
If $\varTheta$ is an infinite cardinal and $\mathcal{E}\prec H(\varTheta)$, then $\mathcal{E}$ contains the finite subsets of itself.
Fact 2.2 (implicit in [soukup2011elementary, Corollary 2.6]).
Let $\aleph_{0}\leq\lambda<\kappa<\varTheta$ be cardinals where $\varTheta$ is regular and let $x\in H(\varTheta)$. Then there is a $\subseteq$-increasing continuous sequence $\left\langle\mathcal{E}_{\alpha}:\ \alpha<\kappa\right\rangle$ where $\mathcal{E}_{0}=\emptyset$ and $\mathcal{E}_{\alpha}\prec H(\varTheta)$ for $1\leq\alpha<\kappa$ such that $x\in\mathcal{E}_{1}$ and for every $\alpha<\kappa$: $\left|\mathcal{E}_{\alpha}\right|\cup\{\mathcal{E}_{\alpha}\}\subseteq\mathcal{E}_{\alpha+1}$ and $\left|\mathcal{E}_{\alpha}\right|=\left|\alpha\right|\cdot\lambda$.
Fact 2.3 ([soukup2011elementary, Claim 3.7]).
Let $\varTheta$ be an uncountable cardinal and let $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$. If $X\in\mathcal{E}$ with $X\setminus\mathcal{E}\neq\emptyset$, then $\left|X\right|>\left|\mathcal{E}\right|$ and $\left|X\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$.
Fact 2.4.
Let $\varTheta$ be an uncountable cardinal, $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$ and let $\mathcal{X}\in\mathcal{E}$ be a set of countable sets. If there is an $X\in\mathcal{X}$ with $X\cap\mathcal{E}=\emptyset$, then for any maximal set $\mathcal{Y}\in\mathcal{E}$ of pairwise disjoint elements of $\mathcal{X}$ we have $\left|\mathcal{Y}\right|>\left|\mathcal{E}\right|$ and $\left|\mathcal{Y}\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$.
Proof.
If $\mathcal{Y}\subseteq\mathcal{E}$, then $\bigcup\mathcal{Y}\subseteq\mathcal{E}$ and $\mathcal{Y}\cup\{X\}$ contradicts the maximality of $\mathcal{Y}$. Therefore $\mathcal{Y}\setminus\mathcal{E}\neq\emptyset$ and we are done by applying Fact 2.3. ∎
2.2. Matroid theory
Matroids were introduced by White [white1992matroid] to generalize the concept of linear independence in vector spaces. In this paper we only work with the most basic concept of infinite matroids, namely finitary matroids. A more general concept of infinite matroids was discovered by Higgs [higgs1969matroids] and rediscovered independently by Bruhn et al. [bruhn2013axioms]. For a survey about basic facts about infinite matroids we refer to [nathanhabil].
An ordered pair $M=(E,\mathcal{C})$ is a finitary matroid (defined by its circuits) if $\mathcal{C}$ is a set of nonempty, pairwise $\subseteq$-incomparable finite subsets of $E$ satisfying the following circuit elimination axiom: If $C_{0},C_{1}\in\mathcal{C}$ are distinct and $e\in C_{0}\cap C_{1}$, then there is a $C\in\mathcal{C}$ with $C\subseteq(C_{0}\cup C_{1})\setminus\{e\}$. The elements of $\mathcal{C}$ are called the circuits of $M$.
Fact 2.5 (Strong circuit elimination).
If $M=(E,\mathcal{C})$ is a finitary matroid, $C_{0},C_{1}\in\mathcal{C}$ with $e\in C_{0}\setminus C_{1}$ and $f\in C_{0}\cap C_{1}$, then there is a $C\in\mathcal{C}$ with $e\in C\subseteq(C_{0}\cup C_{1})\setminus\{f\}$.
Let $F$ be any set [^3] In order to simplify some later notation, we do not even assume $F\subseteq E$.. We write $M\setminus F$ for the matroid with $E(M\setminus F):=E\setminus F$ and $\mathcal{C}(M\setminus F):=\{C\in\mathcal{C}:\ C\cap F=\emptyset\}$. Furthermore, $M\!\!\upharpoonright\!\!F$ stands for $M\setminus(E\setminus F)$.
Fact 2.6.
For every set $F$, the minimal elements of the set $\{C\setminus F:\ C\in\mathcal{C}\}\setminus\{\emptyset\}$ give the circuits of a matroid.
We write $M/F$ for the matroid on $E\setminus F$ where $\mathcal{C}(M/F)$ is the set described in Fact 2.6. Matroids of the form $M/F\setminus G$ are called the minors of $M$. The set $\mathcal{C}^{*}$ of cocircuits of $M$ consists of the minimal nonempty subsets $D$ of $E$ for which $\left|D\cap C\right|\neq 1$ for every $C\in\mathcal{C}$.
Fact 2.7.
For every $D\in\mathcal{C}^{*}$ and for every distinct $e,f\in D$, there is a $C\in\mathcal{C}$ with $C\cap D=\{e,f\}$.
Fact 2.8.
The set $\mathcal{C}^{*}(M\setminus F)$ consists of the minimal elements of $\{D\setminus F:\ D\in\mathcal{C}^{*}\}\setminus\{\emptyset\}$ while $\mathcal{C}^{*}(M/F)=\{D\in\mathcal{C}^{*}:\ D\cap F=\emptyset\}$.
If $G=(V,E)$ is a graph, then the edge sets of the cycles of $G$ are the circuits of a finitary matroid on $E$ which is called the cycle matroid of $G$. Its cocircuits are the bonds of $G$. A vector system with the minimal linearly dependent subsets as circuits always forms a finitary matroid. Such matroids (as well as matroids that are isomorphic to such a matroid) are called linear or representable. If the vector system in question is over $\mathbb{F}_{2}$, then it is a finitary binary matroid. The cycle matroid of a graph is a binary matroid.
Fact 2.9 ([duchamp1989characterizations, Theorem 2.3]).
A finitary matroid $M$ is binary iff $\left|C\cap D\right|$ is even for every $C\in\mathcal{C}$ and $D\in\mathcal{C}^{*}$.
2.3. Oriented matroids
Oriented binary matroids were defined first by Minty in [minty1966axiomatic]. These are also called signed matroids (see [oxley1992book, p. 458]). The more general concept of oriented matroids was discovered independently by Bland and Las Vergnas [bland1978orientability] and by Folkman and Lawrence [folkman1978oriented]. General concepts of infinite oriented matroids were introduced by Hochstättler and Kaspar [hoch2015orthogax]. Infinite binary orientable matroids are characterised (under the name of signable matroids) by Bowler and Carmesin in [bowler2018excluded, p. 111]. Affine representable infinite oriented matroids were investigated recently by Delucchi and Knauer [delucchi2024finitary].
A signing of a set is a ±$\pm 1$-valued function defined on the set. A signed set is positive if it is constant $1$ and negative if it is constant $-1$. The domain of a signed set $X$ is denoted by $\underline{X}$. We define $X^{+}:=\{e\in\underline{X}:\ X(e)=1\}$ and $X^{-}:=\{e\in\underline{X}:\ X(e)=-1\}$. If $\mathcal{X}$ is a set of signed sets, then let $\underline{\mathcal{X}}:=\{\underline{X}:\ X\in\mathcal{X}\}$. A signature of a set $\mathcal{F}$ of nonempty sets is a set that contains exactly two signings of each $F\in\mathcal{F}$ and these two are opposite. An orientation $\vec{M}$ of a finitary matroid $M=(E,\mathcal{C})$ is a pair $(E,\mathcal{O})$ where $\mathcal{O}$ is a signature of $\mathcal{C}$ that satisfies the following oriented circuit elimination axiom: For every $C_{0},C_{1}\in\mathcal{O}$ with $C_{1}\neq\pm C_{0}$ and $e\in C_{0}^{+}\cap C_{1}^{-}$, there is a $C\in\mathcal{C}$ with $C^{-}\subseteq(C_{0}^{-}\cup C_{1}^{-})\setminus\{e\}$ and $C^{+}\subseteq(C_{0}^{+}\cup C_{1}^{+})\setminus\{e\}$. Every oriented matroid gives rise to a unique signature $\mathcal{O}^{*}$ of its cocircuits satisfying
(1) $(\forall C\in\mathcal{O})(\forall D\in\mathcal{O}^{*})\left[\underline{C}\cap\underline{D}\neq\emptyset\Longrightarrow C\!\!\upharpoonright\!\!(\underline{C}\cap\underline{D})\neq\pm D\!\!\upharpoonright\!\!(\underline{C}\cap\underline{D})\right].$
Conversely, if $\mathcal{O}$ and $\mathcal{O}^{*}$ are signatures of the circuits and of the cocircuits respectively and they satisfy orthogonality axiom (1), then $\mathcal{O}$ satisfies the oriented circuit elimination axiom.
Fact 2.10.
A finitary oriented matroid is binary iff $\sum_{e\in\underline{C}\cap\underline{D}}C(e)D(e)=0$ for every $C\in\mathcal{O}$ and $D\in\mathcal{O}^{*}$.
The fact above is well-known in the finite case [oxley1992book, Corollary 13.4.6] and remains true by compactness for finitary matroids.
We refer to the positive and negative (co)circuits of $\vec{M}$ as its directed (co)circuits.
Fact 2.11 (Farkas lemma [bjorner1999oriented, Corollary 3.4.6]).
If $\vec{M}=(E,\mathcal{O})$ is a finite oriented matroid, then for every $e\in E$ there is either a positive $C\in\mathcal{O}$ through $e$ or a positive $D\in\mathcal{O}^{*}$ through $e$ but not both.
Minors of finitary oriented matroids are defined similarly to the unoriented ones, except they inherit the signs as well.
Fact 2.12 ([hoch2015orthogax, Lemma 3.6]).
Let $\vec{M}=(E,\mathcal{O})$ be a finitary oriented matroid and let $F$ be a set. Then
$\mathcal{O}(\vec{M}\setminus F)$ =$=\{C\in\mathcal{O}:\ \underline{C}\cap F=\emptyset\},$
$\mathcal{O}(\vec{M}/F)$ =$=\{C\!\!\upharpoonright\!\!(E\setminus F):\ C\in\mathcal{O},\ \underline{C}\setminus F\in\mathcal{C}(M/F)\},$
$\mathcal{O}^{*}(\vec{M}\setminus F)$ =$=\{D\!\!\upharpoonright\!\!(E\setminus F):\ D\in\mathcal{O}^{*},\ \underline{D}\setminus F\in\mathcal{C}^{*}(M/F)\},$
$\mathcal{O}^{*}(\vec{M}/F)$ =$=\{D\in\mathcal{O}^{*}:\ \underline{D}\cap F=\emptyset\}.$
3. Finite matching extendable matroids
Using the terminology of Komjáth et al. from [komjath1988compactness], a matching in matroid $M=(E,\mathcal{C})$ is a set $\mathcal{A}$ of pairwise disjoint elements of $\mathcal{C}$. A perfect matching is a matching $\mathcal{A}$ with $\bigcup\mathcal{A}=E$, i. e. a partition of the ground set of $M$ into circuits. A matroid is matchable if it admits a perfect matching. A matroid $M$ is called finite matching extendable (shortly f.m.e.) if for every finite matching $\mathcal{A}$ and for every edge $e\in E\setminus\bigcup\mathcal{A}$ there is a circuit $C$ with $e\in C\subseteq E\setminus\bigcup\mathcal{A}$.
Theorem 3.1 (Komjáth, Milner, Polat; [komjath1988compactness, Theorem 1]).
Every finitary f.m.e. matroid is matchable.
Proof.
Let $\varTheta$ be an arbitrary but fixed regular uncountable cardinal. It is sufficient to prove the theorem only for matroids in $H(\varTheta)$.
Observation 3.2.
If $\mathcal{E}\prec H(\varTheta)$, then for every finitary f.m.e. matroid $M\in\mathcal{E}$, the matroid $M\!\!\upharpoonright\!\!\mathcal{E}$ is also f.m.e.
Proof.
Let a finite matching $\mathcal{A}$ in $M\!\!\upharpoonright\!\!\mathcal{E}$ and an edge $e\in E\cap\mathcal{E}\setminus\bigcup\mathcal{A}$ be given. Then for each $C\in\mathcal{A}$ we have $C\in\mathcal{E}$, moreover, $\mathcal{A}\in\mathcal{E}$ (see Fact 2.1). Thus $M,\mathcal{A},e\in\mathcal{E}$ and hence there is a $C\in\mathcal{C}\cap\mathcal{E}$ with $e\in C\subseteq E\setminus\bigcup\mathcal{A}$. Finally, $C\subseteq\mathcal{E}$ because $C$ is finite (see Fact 2.3), therefore $C\in\mathcal{C}(M\!\!\upharpoonright\!\!\mathcal{E})$. ∎
Lemma 3.3.
If $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$, then for every finitary f.m.e. matroid $M\in\mathcal{E}$, the matroid $M\setminus\mathcal{E}$ is also f.m.e.
Proof.
Let a finite matching $\mathcal{A}$ in $M\setminus\mathcal{E}$ and an edge $e\in E\setminus(\bigcup\mathcal{A}\cup\mathcal{E})$ be given. By applying the assumption that $M$ is f.m.e., we pick a $C_{0}\in\mathcal{C}$ with $e\in C_{0}\subseteq E\setminus\bigcup\mathcal{A}$ for which $F:=C_{0}\cap\mathcal{E}$ is minimal. If $F=\emptyset$, then $C_{0}\subseteq E\setminus\mathcal{E}$ and hence $C_{0}\in\mathcal{C}(M\setminus\mathcal{E})$, thus we are done. Suppose for a contradiction that $F\neq\emptyset$. Let $\mathcal{X}$ consist of the finite sets that extend $F$ to an $M$-circuit, formally $\mathcal{X}:=\{X\subseteq E\setminus F:\ F\cup X\in\mathcal{C}\}$. Since $C_{0}\setminus\mathcal{E}\in\mathcal{X}$ is clearly disjoint from $\mathcal{E}$, we conclude by applying Fact 2.4 that there is a set $\mathcal{Y}$ of pairwise disjoint elements of $\mathcal{X}$ with $\left|\mathcal{Y}\right|>\left|\mathcal{E}\right|$. Since $\left|\mathcal{E}\cup\bigcup\mathcal{A}\cup\{e\}\right|=\left|\mathcal{E}\right|$, it follows that there is an $X\in\mathcal{Y}$ that is disjoint from $\mathcal{E}\cup\bigcup\mathcal{A}\cup\{e\}$. Let $C_{1}:=F\cup X$. Since $F\neq\emptyset$ by the indirect assumption, we can pick $f\in C_{0}\cap C_{1}$. By strong circuit elimination (Fact 2.5), it follows that there is a $C\in\mathcal{C}$ with $e\in C\subseteq(C_{0}\cup C_{1})\setminus\{f\}$. But then $e\in C\subseteq E\setminus\bigcup\mathcal{A}$ with $\left|C\cap\mathcal{E}\right|\leq\left|C_{0}\cap\mathcal{E}\right|-1$ (because $C\cap\mathcal{E}\subseteq C_{0}\cap\mathcal{E}\setminus\{f\}$) which contradicts the choice of $C_{0}$. ∎
Let $M=(E,\mathcal{C})\in H(\varTheta)$ be a f.m.e. matroid. We apply transfinite induction on $\kappa:=\left|E\right|$. If $\kappa\leq\aleph_{0}$, the statement follows by a straightforward recursion. Indeed, we fix a $\kappa$-type well-order of $E$ and in step $n$ apply finite matching extendability to pick a circuit $C_{n}$ through the smallest element of $E\setminus\bigcup_{i<n}C_{i}$ that is disjoint from $C_{0},\dots,C_{n-1}$.
Suppose $\kappa>\aleph_{0}$. Let $\left\langle\mathcal{E}_{\alpha}:\ \alpha<\kappa\right\rangle$ be as in Fact 2.2 with $x:=M$ and $\lambda:=\aleph_{0}$. By Lemma 3.3 we know that $M\setminus\mathcal{E}_{\alpha}$ is f.m.e. for every $\alpha<\kappa$. Since $M,\mathcal{E}_{\alpha}\in\mathcal{E}_{\alpha+1}$, we have $M\setminus\mathcal{E}_{\alpha}\in\mathcal{E}_{\alpha+1}$. But then applying Observation 3.2 with $M\setminus\mathcal{E}_{\alpha}$ and $\mathcal{E}_{\alpha+1}$ for $\alpha<\kappa$, we conclude that $(M\setminus\mathcal{E}_{\alpha})\!\!\upharpoonright\!\!\mathcal{E}_{\alpha+1}=M\!\!\upharpoonright\!\!(\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha})$ is f.m.e. for every $\alpha<\kappa$. By the induction hypothesis we can pick a perfect matching $\mathcal{A}_{\alpha}$ of $M\!\!\upharpoonright\!\!(\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha})$ for each $\alpha<\kappa$. Clearly, $\bigcup_{\alpha<\kappa}\mathcal{A}_{\alpha}$ is a matching in $M$ and it is perfect because $E=\bigcup_{\alpha<\kappa}E\cap(\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha})$ is a partition. ∎
4. A Laviolette theorem for finitary matroids
Let $M=(E,\mathcal{C})$ be a finitary matroid and let $\lambda$ be an infinite cardinal. An $F\subseteq E$ is $\lambda$-cocircuit faithful w.r.t. $M$ if for every $D\in\mathcal{C}^{*}(M\!\!\upharpoonright\!\!F)$ with $\left|D\right|<\lambda$ we have $D\in\mathcal{C}^{*}(M)$. A $\lambda$-cocircuit faithful partition of $M$ is a partition $\{F_{i}:\ i\in I\}$ of $E$ such that
  1. (i)
    $\left|F_{i}\right|\leq\lambda$ for every $i\in I$;
  2. (ii)
    $F_{i}$ is $\lambda$-cocircuit faithful for each $i\in I$;
  3. (iii)
    For every $D\in\mathcal{C}^{*}$ with $\left|D\right|\leq\lambda$, there exists an $i\in I$ with $D\subseteq F_{i}$.
Theorem 4.1 (Laviolette theorem for finitary matroids).
For every infinite cardinal $\lambda$, every finitary matroid $M=(E,\mathcal{C})$ admits a $\lambda$-cocircuit faithful partition.
Proof.
Let $\varTheta$ be an arbitrary but fixed regular uncountable cardinal. It is sufficient to prove the theorem only for matroids in $H(\varTheta)$.
Lemma 4.2.
Let $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$. Then for every finitary matroid $M\in\mathcal{E}$ and for every $D\in\mathcal{C}^{*}$ with $D\cap\mathcal{E}\neq\emptyset$ we have: $D\subseteq\mathcal{E}$ if $\left|D\right|\leq\left|\mathcal{E}\right|$ and $\left|D\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$ if $\left|D\right|>\left|\mathcal{E}\right|$.
Proof.
We may assume that $D\setminus\mathcal{E}\neq\emptyset$ since if $D\subseteq\mathcal{E}$, then we are done. We need to show that $\left|D\right|>\left|\mathcal{E}\right|$ and $\left|D\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$. Let $e\in D\cap\mathcal{E}$ and $f\in D\setminus\mathcal{E}$. By Fact 2.7, there is a $C\in\mathcal{C}$ with $C\cap D=\{e,f\}$. Let $F:=C\cap\mathcal{E}$ and $\mathcal{X}:=\{X:\ F\cup X\in\mathcal{C}\}$. Note that each $X\in\mathcal{X}$ meets $D$ since otherwise $(F\cup X)\cap D=\{e\}$ contradicts $D\in\mathcal{C}^{*}$. Clearly, $C\setminus\mathcal{E}\in\mathcal{X}$ is disjoint from $\mathcal{E}$. Let $\mathcal{Y}\in\mathcal{E}$ be a maximal family of pairwise disjoint elements of $\mathcal{X}$. By Fact 2.4 we know that $\left|\mathcal{Y}\right|>\left|\mathcal{E}\right|$ and $\left|\mathcal{Y}\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$. Since the elements of $\mathcal{Y}$ are pairwise disjoint and each of them meets $D$, we conclude that $\left|D\right|>\left|\mathcal{E}\right|$ and $\left|D\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$. ∎
Let $M=(E,\mathcal{C})\in H(\varTheta)$ be a finitary matroid. We apply transfinite induction on $\kappa:=\left|E\right|$. If $\kappa\leq\lambda$, then the trivial partition $\{E\}$ is suitable. If $\kappa>\lambda$, then let $\left\langle\mathcal{E}_{\alpha}:\ \alpha<\kappa\right\rangle$ be as in Fact 2.2 where $x=M$. By Lemma 4.2, for every $D\in\mathcal{C}^{*}$ and for the smallest ordinal $\alpha+1$ for which $D\cap\mathcal{E}_{\alpha+1}\neq\emptyset$, we have $D\subseteq\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha}$ if $\left|D\right|\leq\lambda$ and $\left|D\cap(\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha})\right|=\lambda$ if $\left|D\right|>\lambda$. We define $F_{\alpha}:=E\cap(\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha})$ for $\alpha<\lambda$. Then the partition $E=\bigcup_{\alpha<\lambda}F_{\alpha}$ satisfies (ii) and (iii), moreover $\left|F_{\alpha}\right|<\kappa$ for each $\alpha<\kappa$. Apply the induction hypothesis to $M\!\!\upharpoonright\!\!F_{\alpha}$ and let $F_{\alpha}=\bigcup_{i\in I_{\alpha}}F_{\alpha,i}$ be the resulting partition. We claim that $E=\bigcup_{\alpha<\lambda,i\in I_{\alpha}}F_{\alpha,i}$ is a $\lambda$-cocircuit faithful partition of $M$. Property (i) is clear by the induction hypothesis. If $D\in\mathcal{C}^{*}(M\!\!\upharpoonright\!\!F_{\alpha,i})$ with $\left|D\right|<\kappa$, then by induction $D\in\mathcal{C}^{*}(M\!\!\upharpoonright\!\!F_{\alpha})$. But then $D\in\mathcal{C}^{*}$ because partition $E=\bigcup_{\alpha<\lambda}F_{\alpha}$ satisfies (ii). Therefore (ii) holds true. Finally, let $D\in\mathcal{C}^{*}(M)$ with $\left|D\right|\leq\kappa$. Then there is an $\alpha$ with $D\in\mathcal{C}^{*}(M\!\!\upharpoonright\!\!F_{\alpha})$ because partition $E=\bigcup_{\alpha<\lambda}F_{\alpha}$ satisfies (iii). But then by induction there is an $i\in I_{\alpha}$ such that $D\in\mathcal{C}^{*}(M\!\!\upharpoonright\!\!F_{\alpha,i})$. Thus (iii) holds as well. ∎
5. Farkas lemma in finitary oriented matroids
First, we give a simple example due to Kaspar showing that the Farkas lemma (Fact 2.11) does not extend to finitary oriented matroids. Then we prove that it does extend for binary ones.
Claim 5.1.
There exists a finitary oriented matroid $\vec{M}=(E,\mathcal{O})$ that has neither a positive circuit nor a positive cocircuit.
Proof.
Let $\vec{M}:=(\mathbb{Z},\mathcal{O})$ where $\underline{\mathcal{O}}$ consists of the $3$-element subsets of $\mathbb{Z}$ and for $i<j<k$ the signed set $\{\left\langle i,1\right\rangle,\left\langle j,-1\right\rangle,\left\langle k,1\right\rangle\}$ and its negation are in $\mathcal{O}$. Then $\underline{\mathcal{O}^{*}}=\{\mathbb{Z}\setminus\{n\}:\ n\in\mathbb{Z}\}$ and $\mathcal{O}^{*}$ consists of the signed set $\{\left\langle m,1\right\rangle:\ m<n\}\cup\{\left\langle m,-1\right\rangle:\ n<m\}$ and its negation for each $n\in\mathbb{Z}$. It is routine to check that the orthogonality axiom 1 holds, thus $\mathcal{O}$ and $\mathcal{O}^{*}$ are indeed the circuit signature and cocircuit signature of a matroid respectively. ∎
Lemma 5.2 (Farkas lemma for finitary binary oriented matroids).
Let $\vec{M}=(E,\mathcal{O})$ be a finitary binary orientated matroid. Then for every $e\in E$, there is either a positive circuit or a positive cocircuit through $e$ but not both.
Proof.
We cannot have both by Fact 2.10. Suppose for a contradiction that there is neither a positive circuit nor a positive cocircuit through $e_{0}\in E$ in $\vec{M}$. By Fact 2.11 and by the non-existence of positive circuits through $e_{0}$, we know that for every finite $E^{\prime}\subseteq E$ containing $e_{0}$, we can find a positive cocircuit of $\vec{M}\!\!\upharpoonright\!\!E^{\prime}$ through $e_{0}$. Thus by compactness there exists a set $F\subseteq E$ containing $e_{0}$ such that $\sum_{e\in F\cap\underline{C}}C(e)=0$ for every $C\in\mathcal{C}$. By Zorn’s lemma we can assume that $F$ is minimal w.r.t. these properties. Clearly, $\left|F\cap\underline{C}\right|\neq 1$ for each $C\in\mathcal{C}$, thus by Zorn’s lemma $F$ includes a cocircuit of $M$ through $e_{0}$. Thus there is a $D\in\mathcal{O}^{*}$ with $e_{0}\in\underline{D}\subseteq F$. By negating $D$ if necessary we can assume $D(e_{0})=1$. By the indirect assumption $D$ is not a positive cocircuit, thus there is an $e_{1}\in\underline{D}$ with $D(e_{1})=-1$. Let $\mymathbb{1}_{F}$ be the characteristic function of $F$ (defined on $E$) and let $\widehat{D}$ be the extensions of $D$ to the domain $E$ with 00 values. We define $f:=\widehat{D}+\mymathbb{1}_{F}$. Then
  1. (i)
    $f\geq 0$,
  2. (ii)
    $f(e_{0})=2$,
  3. (iii)
    $\sum_{e\in\underline{C}}f(e)C(e)=0$ for every $C\in\mathcal{O}$.
For the support $S:=\{e\in E:\ f(e)\neq 0\}$ of $f$ we have $S\subseteq F\setminus\{e_{1}\}\subsetneq F$. Thus in order to get a contradiction to the minimality of $F$, we will construct by compactness a $G$ with $e_{0}\in G\subseteq S$ such that $\sum_{e\in G\cap\underline{C}}C(e)=0$ for every $C\in\mathcal{O}$.
Let $S^{\prime}\subseteq S$ be any finite set containing $e_{0}$. If $(\vec{M}.S)\!\!\upharpoonright\!\!S^{\prime}$ has a positive circuit $C$ through $e_{0}$, then by Fact 2.12 there is a $C^{\prime}\in\mathcal{O}$ with $C^{\prime}\!\!\upharpoonright\!\!S=C$. But then by applying (i) and (ii) together with the positivity of $C$ we obtain
$\sum_{e\in\underline{C^{\prime}}}C^{\prime}(e)f(e)=\sum_{e\in\underline{C}}C(e)f(e)\geq f(e_{0})C(e_{0})=2,$
which contradicts (iii). Hence there is no such a circuit and therefore $(M.S)\!\!\upharpoonright\!\!S^{\prime}$ has a positive cocircuit $D^{\prime}$ through $e_{0}$ by Fact 2.11. Then $e_{0}\in\underline{D^{\prime}}\subseteq S$ and for every $C\in\mathcal{O}$ with $\underline{C}\subseteq S^{\prime}$ we have $\sum_{e\in\underline{D^{\prime}}\cap\underline{C}}C(e)=0$. Since the finite set $S^{\prime}$ was arbitrary, it follows by compactness that the desired $G$ exists, which concludes the proof. ∎
6. Partitioning oriented matroids into directed circuits
In an oriented matroid $\vec{M}$ a $D\in\mathcal{O}^{*}$ is called balanced if $\left|D^{+}\right|=\left|D^{-}\right|$. We call $\vec{M}$ itself balanced if every $D\in\mathcal{O}^{*}$ is balanced.
Theorem 6.1.
A finitary binary oriented matroid can be partitioned into directed circuits iff it is balanced.
Proof.
Let $\vec{M}=(E,\mathcal{O})$ be a finitary oriented matroid. Suppose that $\mathcal{A}$ is a set of directed circuits, say positive ones, such that $\underline{\mathcal{A}}$ is a partition of $E$. Then for any $D\in\mathcal{O}^{*}$ and $s\in\{+,-\}$, we have $D^{s}=\dot{\bigcup}\{\underline{C}\cap D^{s}:\ C\in\mathcal{A}\}$. It follows from Fact 2.10, that we must have $\left|\underline{C}\cap D^{+}\right|=\left|\underline{C}\cap D^{-}\right|$ for every positive circuit $C$, therefore
$\left|D^{+}\right|=\sum_{C\in\mathcal{A}}\left|\underline{C}\cap D^{+}\right|=\sum_{C\in\mathcal{A}}\left|\underline{C}\cap D^{-}\right|=\left|D^{-}\right|.$
We turn to the non-trivial direction. Let a regular uncountable cardinal $\varTheta$ be fixed. We need a couple of lemmas:
Lemma 6.2.
If $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$, then for every finitary balanced oriented matroid $\vec{M}\in\mathcal{E}$, the matroid $\vec{M}\setminus\mathcal{E}$ is also balanced.
Proof.
Let $D_{0}\in\mathcal{O}^{*}(\vec{M}\setminus\mathcal{E})$ be given. By Fact 2.12 there is a $D_{1}\in\mathcal{O}^{*}(\vec{M})$ such that $D_{0}=D_{1}\!\!\upharpoonright\!\!(E\setminus\mathcal{E})$. If $\underline{D_{1}}\cap\mathcal{E}=\emptyset$, then $D_{0}=D_{1}\in\mathcal{O}^{*}(\vec{M})$ and we are done because $D_{1}$ is balanced by assumption. Suppose that $D_{1}\cap\mathcal{E}\neq\emptyset$. Then it follows by applying Lemma 4.2 with $\underline{D_{1}}$ and $M$ that we must have $\left|\underline{D_{1}}\right|>\left|\mathcal{E}\right|$. Since $\left|D_{1}\setminus D_{0}\right|<\left|\mathcal{E}\right|$, we conclude that $\left|D_{1}^{s}\right|=\left|D_{0}^{s}\right|$ for $s\in\{+,-\}$, therefore
$\left|D_{0}^{+}\right|=\left|D_{1}^{+}\right|=\left|D_{1}^{-}\right|=\left|D_{0}^{-}\right|.$
Lemma 6.3.
If $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$, then for every finitary binary balanced oriented matroid $\vec{M}\in\mathcal{E}$, the matroid $\vec{M}\!\!\upharpoonright\!\!\mathcal{E}$ is also balanced.
Proof.
Let $D\in\mathcal{O}^{*}(\vec{M}\!\!\upharpoonright\!\!\mathcal{E})$ be arbitrary. By symmetry we may assume that $\left|D^{+}\right|\geq\left|D^{-}\right|$. We need to show that $\left|D^{+}\right|\leq\left|D^{-}\right|$. Clearly, $D^{+}\neq\emptyset$ because $D\neq\emptyset$. Fix an $e_{0}\in D^{+}$ and let $\vec{N}:=\vec{M}/(\mathcal{E}\setminus\underline{D})\setminus(\underline{D}\setminus\{e_{0}\})$. We apply Lemma 5.2 with $e_{0}$ and $\vec{N}$. Suppose first that there is a positive $D_{0}\in\mathcal{O}^{*}(\vec{N})$ through $e_{0}$. Then there is a $D_{1}\in\mathcal{O}^{*}(\vec{M})$ with $D_{1}\!\!\upharpoonright\!\!((E\setminus\underline{D})\cup\{e_{0}\})=D_{0}$ (see Fact 2.12). Then $e_{0}\in\underline{D_{1}}\cap\mathcal{E}\subseteq\underline{D}$ by construction and hence we must have $\underline{D_{1}}\cap\mathcal{E}=\underline{D}$ because $\underline{D_{1}}\cap\mathcal{E}$ includes a cocircuit of $\vec{M}\!\!\upharpoonright\!\!\mathcal{E}$ by Fact 2.8 but this cannot be a proper subset of another cocircuit $D$. Then $D_{1}\!\!\upharpoonright\!\!\mathcal{E}=\pm D$ and by $D_{1}(e_{0})=D(e_{0})=1$ we conclude that $D_{1}\!\!\upharpoonright\!\!\mathcal{E}=D$. But then by the positivity of $D_{1}\setminus D$ and by the assumption that $\vec{M}$ is balanced:
$\left|D^{+}\right|\leq\left|D_{1}^{+}\right|=\left|D_{1}^{-}\right|=\left|D^{-}\right|.$
Assume now that there is a positive $C_{0}\in\mathcal{O}(\vec{N})$ through $e_{0}$. Then there is a $C_{1}\in\mathcal{O}(\vec{M})$ with $C_{1}\!\!\upharpoonright\!\!(E\setminus(\mathcal{E}\setminus\underline{D}))=C_{0}$. Let $F:=C_{1}\cap\mathcal{E}$ and let $\mathcal{X}\in\mathcal{E}$ be the set of those positive signed sets $X$ for which $F\cup X\in\mathcal{O}(\vec{M})$. Then $\underline{\mathcal{X}}$ contains an element, namely $\underline{C_{1}}\setminus\mathcal{E}$, that is disjoint from $\mathcal{E}$. It follows by Fact 2.4 that one can find $\left|\mathcal{E}\right|$ many pairwise disjoint elements in $\underline{\mathcal{X}}\cap\mathcal{E}$. To prove $\left|D^{+}\right|\leq\left|D^{-}\right|$, it is enough to show that for each $X\in\mathcal{X}\cap\mathcal{E}$, there is an $e\in\underline{D}\cap\underline{X}$ with $D(e)=-1$. Suppose that $X$ is a counterexample, and let $D^{\prime}\in\mathcal{O}^{*}(\vec{M})$ with $D^{\prime}\!\!\upharpoonright\!\!\mathcal{E}=D$. Note that $\underline{D^{\prime}}\cap(\underline{F}\cup\underline{X})=\underline{D}\cap(\underline{F}\cup\underline{X})$ because $\underline{F}\cup\underline{X}\subseteq\mathcal{E}$ and recall that $\underline{F}\cap\underline{D}=\{e_{0}\}$ by construction. Then by using the indirect assumption and the positivity of $X$:
$1=D(e_{0})F(e_{0})\leq\sum_{e\in\underline{D}\cap(\underline{X}\cup\underline{F})}D(e)(F\cup X)(e)=\sum_{e\in\underline{D^{\prime}}\cap(\underline{X}\cup\underline{F})}D^{\prime}(e)(F\cup X)(e).$
But $D^{\prime}\in\mathcal{O}^{*}$ and $F\cup X\in\mathcal{O}$, thus this contradicts Fact 2.10. ∎
We prove by induction on $\kappa:=\left|E\right|$ that every finitary binary balanced oriented matroid $\vec{M}=(E,\mathcal{O})\in H(\varTheta)$ admits a partition into positive circuits.
Observation 6.4.
In a finitary binary balanced oriented matroid every edge is contained in a positive circuit.
Proof.
If follows directly from Lemma 2.11 because being balanced ensures the non-existence of positive cocircuits. ∎
Observation 6.5.
If $\mathcal{A}$ is a finite set of pairwise disjoint positive circuits of $\vec{M}$, then $\vec{M}\setminus\bigcup\underline{\mathcal{A}}$ is balanced.
Proof.
Let $D\in\mathcal{O}^{*}(\vec{M}\setminus\bigcup\underline{\mathcal{A}})$. Pick a $D^{\prime}\in\mathcal{O}^{*}$ with $D^{\prime}\!\!\upharpoonright\!\!(E\setminus\bigcup\underline{\mathcal{A}})=D$. Since $\vec{M}$ is binary and each $C\in\mathcal{A}$ is positive, Fact 2.10 ensures that $D^{\prime}$ has the same number of positive and negative edges in $\bigcup\underline{\mathcal{A}}$. But then so does $D$ because $D^{\prime}$ is balanced by assumption. ∎
If $\kappa\leq\aleph_{0}$, then the desired partition can be obtained by a straightforward recursion via Observations 6.4 and 6.5. Indeed, we fix a $\kappa$-type well-order of $E$ and in step $n$ apply Observation 6.4 to $\vec{M}\setminus\bigcup_{i<n}\underline{C_{i}}$ (which is balanced by Observations 6.5) to find a positive circuit $C_{n}$ disjoint from $C_{0},\dots,C_{n-1}$ through the smallest element of $E\setminus\bigcup_{i<n}C_{i}$
Suppose $\kappa>\aleph_{0}$. Let $\left\langle\mathcal{E}_{\alpha}:\ \alpha<\kappa\right\rangle$ be as in Fact 2.2 with $x=\vec{M}$ and $\lambda=\aleph_{0}$. Then for each $\alpha<\kappa$: $\vec{M}\setminus\mathcal{E}_{\alpha}$ is balanced by Lemma 6.2. Since $\vec{M},\mathcal{E}_{\alpha}\in\mathcal{E}_{\alpha+1}$, we have $\vec{M}\setminus\mathcal{E}_{\alpha}\in\mathcal{E}_{\alpha+1}$, thus by Lemma 6.3 we conclude that $(\vec{M}\setminus\mathcal{E}_{\alpha})\!\!\upharpoonright\!\!\mathcal{E}_{\alpha+1}$ is balanced. By the induction hypotheses, there is a partition $\mathcal{A}_{\alpha}$ of $(\vec{M}\setminus\mathcal{E}_{\alpha})\!\!\upharpoonright\!\!\mathcal{E}_{\alpha+1}$ into positive circuits. Finally, $\bigcup_{\alpha<\kappa}\mathcal{A}_{\alpha}$ is a partition of $\vec{M}$ into positive circuits. ∎
[…]