### v2

# Circuit-partition of infinite matroids

Abstract.
Komjáth, Milner, and Polat investigated when a finitary matroid admits a partition into circuits. They defined the class of “finite matching extendable” matroids and showed in their compactness theorem that those matroids always admit such a partition. Their proof is based on Shelah’s singular compactness technique and a careful analysis of certain $\triangle$-systems.
We provide a short, simple proof of their theorem. Then we show that a finitary binary oriented matroid can be partitioned into directed circuits if and only if, in every cocircuit, the cardinality of the negative and positive edges is the same. This generalizes an earlier conjecture of Thomassen, settled affirmatively by the second author, about partitioning the edges of an infinite directed graph into directed cycles. As side results, a Laviolette theorem for finitary matroids and a Farkas lemma for finitary binary oriented matroids are proven. An example is given to show that, in contrast to finite oriented matroids, ‘binary’ is essential in the latter result.
Key words and phrases:
Finitary matroid, Circuit partition, Farkas lemma, Laviolette’s theorem
2020 Mathematics Subject Classification:
Primary: 05B35 Secondary: 03E05, 05A18, 05C63
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Grant No. 513023562 and partially by NKFIH OTKA-129211
1. Introduction
Decomposing complex structures into simple parts is a common theme in mathematical research, extending beyond combinatorics. Cycles are among the most fundamental graph classes, leading to a natural question: When is it possible to partition a graph into cycles? For edge-partitions of finite graphs, the answer is straightforward. On the one hand, each cycle contributes an even amount to the degree of each vertex, so the exclusion of vertices of odd degree is necessary. On the other hand, if all degrees are even, a desired edge-partition can be constructed “greedily”. The problem becomes significantly harder for infinite graphs, especially uncountable ones. The absence of vertices of odd degree remains necessary but is not sufficient. For example, in a two-way infinite path, each vertex has degree two, yet it does not contain a single cycle. Nash-Williams proved in his seminal paper [19, p. 235 Theorem 3] that excluding odd cuts [^1] Finite cuts with an odd number of edges. is sufficient (and obviously necessary). New, simpler proofs of his theorem were found by L. Soukup [23, Theorem 5.1] and Thomassen [24]. In the same paper, Thomassen conjectured that the edges of a directed graph can be partitioned into directed cycles if and only if, in every cut, the cardinality of edges going in one direction equals the cardinality of edges going in the opposite direction. This was confirmed by the second author [15, Theorem 2].
Welsh observed [25] that a finite binary matroid [^2] Definitions and basic facts about finitary matroids are given in Subsection 2.2. can be partitioned into circuits if and only if it has no odd cocircuit, which generalizes the analogous observation about graphs. Polat extended this result to finitary binary matroids [22, Theorem 2.1 (iv) and (vi)], providing a common generalization of Nash-Williams’ theorem and Welsh’s result. Later, together with Komjáth and Milner [16, Theorem 1], he further generalized his result by proving that matroids, they call “finite matching extendable” (f.m.e.), can always be partitioned into circuits (they call matroids with this property “matchable”): {restatable*}[Komjáth, Milner, Polat; [16, Theorem1]]theoremfme Every finitary f.m.e. matroid is matchable.
Since binary matroids that do not contain odd cocircuits are f.m.e., this is indeed a generalization. Their proof is based on Shelah’s singular compactness technique and a careful analysis of certain $\triangle$-systems. We begin by providing a short and relatively simple proof of their theorem (Theorem 2).
A natural question is whether it is also possible to generalize the directed cycle edge-partition result [15, Theorem 2] to finitary binary oriented matroids, as Polat generalized Nash-Williams’ theorem to finitary binary matroids. The first difficulty corresponds to the so-called Farkas lemma. In a directed graph, it is easy to see that every directed edge is contained in a directed cycle or a directed cut. This is known to be true for finite oriented matroids, but we will show that it may fail severely in infinite ones (see Claim 5.1). However, we prove that it holds true in the binary case: {restatable*}[Farkas lemma for finitary binary oriented matroids]lemmaFarinf Let $\vec{M}=(E,\mathcal{O})$ be a finitary binary oriented matroid. Then for every $e\in E$, there is either a positive circuit or a positive cocircuit through $e$ but not both.
Using this, we prove that a finitary binary oriented matroid can be partitioned into directed circuits if and only if, in every signed cocircuit, the cardinality of the negative and positive edges is the same (in which case we refer to that oriented matroid as “balanced”): {restatable*}theoremorparti A finitary binary oriented matroid can be partitioned into directed circuits if and only if it is balanced.
Let $\lambda$ be an infinite cardinal. An edge-partition $\{G_{i}:i\in I\}$ of a graph $G$ into subgraphs of size at most $\lambda$ is called $\lambda$-bond faithful if in each $G_{i}$ every bond of size less than $\lambda$ is also a bond of $G$, and each bond of $G$ of size at most $\lambda$ is a bond of a suitable $G_{i}$. Laviolette proved that an $\aleph_{0}$-bond faithful partition always exists, and, under the Generalized Continuum Hypothesis, $\lambda$-bond faithful partitions exist for every $\lambda$ (see [17, Proposition 3]). L. Soukup showed that the Generalized Continuum Hypothesis can be omitted (see [23, Theorem 6.2]). We define $\lambda$-cocircuit faithful partitions of matroids analogously to $\lambda$-bond faithful partitions, and prove the following generalization of Laviolette’s theorem: {restatable*}[Laviolette theorem for finitary matroids]theoremLaviolette For every infinite cardinal $\lambda$, every finitary matroid $M=(E,\mathcal{C})$ admits a $\lambda$-cocircuit faithful partition.
This paper is organized as follows. In the next section, we survey the basic facts and notation related to set theory and matroid theory that we will use. Sections 3 and 4, due to the second author, discuss the new short proof of the compactness theorem by Komjáth et al. and the generalization of Laviolette’s theorem. Sections 5 and 6 are based on previous works of the first author with Carmesin, who kindly gave us permission to use them. Their respective contents are the generalization of the Farkas lemma and the directed circuit partition result. The framework of our proofs involves cutting up uncountable matroids by a chain of elementary submodels, a method popularized by L. Soukup in [23]. Finally, we raise a couple of open questions in Section 7.
2. Notation and preliminaries
2.1. Set theory
We use variables $\alpha$ and $\beta$ to stand for ordinal numbers, while $\kappa$, $\lambda$ and $\Theta$ denote cardinals. An infinite cardinal $\kappa$ is called regular if it cannot be expressed as the sum of fewer than $\kappa$ many cardinals, each of which is less than $\kappa$. We write $\bigcup X$ for the union of the sets in $X$. A function $f$ is a set of ordered pairs such that if $\left\langle x,y\right\rangle,\left\langle x,z\right\rangle\in f$, then $y=z$. If $X$ is any set, (not necessarily $X\subseteq\mathsf{dom}(f)$), then we write $f\!\!\upharpoonright\!\!X$ for $\{\left\langle x,y\right\rangle\in f:\ x\in X\}$.
2.1.1. Compactness principle
The compactness principle in infinite combinatorics allows us to use finite reasoning to prove infinite results. A little more precisely, suppose that we want to choose suitable values for each of infinitely many variables, each of which can take any one of finitely many possible values. If we must choose these values subject to an infinite set of constraints, each of which depends only on finitely many variables, then the compactness principle says that we can set the variables to satisfy all the constraints if we can do so for every finite subset of the constraints. The first historical application of the compactness principle in infinite combinatorics is a short proof of a famous De Bruijn–Erdős theorem [7] due to Gottschalk (see in [12]). The theorem states that if every finite subgraph of an infinite graph $G$ is $k$-colourable [^3] A graph is $k$-colourable if its vertices can be coloured by at most $k$ colours in such a way that the endpoints of each edge are assigned different colours. for a $k<\omega$, then so is $G$. We give here a brief demonstration of the compactness principle by repeating Gottschalk’s proof.
Recall that a topological space $(X,\tau)$ is called compact if whenever $\mathcal{F}$ is a set of closed subsets of $X$ such that every finite subfamily of $\mathcal{F}$ has nonempty intersection, then so does the whole family $\mathcal{F}$. Consider the compact space consisting of $\{1,\dots,k\}$ equipped with the discrete topology. For $v\in V(G)$, let $(X_{v},\tau_{v})$ be a copy of this space and let $(X,\tau):=\prod_{v\in V(G)}(X_{v},\tau_{v})$. Then $(X,\tau)$ is compact by Tychonoff’s theorem. For $uv\in E(G)$, let $C_{uv}\subseteq X$ consists of those $x$ for which $x(u)\neq x(v)$. Then $\mathcal{F}:=\{C_{e}:\ e\in E(G)\}$ is a family of closed sets. In order to show that $G$ is $k$-colourable, we need to prove that the intersection of all the sets in $\mathcal{F}$ is nonempty. The intersection of any finitely many sets in $\mathcal{F}$ is nonempty because the finite subgraphs of $G$ are properly $k$-colourable by assumption. But then by the compactness of the space $(X,\tau)$, the intersection of all sets in $\mathcal{F}$ is also nonempty.
In this example the variables to be chosen are the colours of the vertices of $G$, and the closure of each $C_{e}$ follows from the fact that the constraint it encodes only mentions finitely many (in this case just two) of these variables.
For a more detailed introduction to the compactness principle in combinatorics, with more examples, see [9].
2.1.2. Elementary submodels
Dividing an uncountable structure into smaller pieces in order to use an inductive argument is a common strategy in proofs about uncountable structures. In order to make the induction hypothesis applicable, one needs to make sure that the pieces inherit certain properties that are relevant with respect to the problem. The desired properties would typically state that these pieces are closed under certain operations. Constructing such closed pieces “by hand” takes a lot of effort and often requires a long description of various operations, tailored to the problem, to close under.
This is where a tool from Model Theory comes in useful, namely elementary submodels. These are a kind of substructure which are, informally speaking, closed under every reasonable operation you could think of. Furthermore, there are standard theorems which allow us to extend arbitrary substructures to elementary ones without increasing their cardinality. Thus the theory of elementary submodels provides a simple, efficient way to make the pieces closed under all possible operations without explicitly describing any of those operations.
A typical approach is to pick a set that is big enough to be a “good approximation” of the whole set-theoretic universe [^4] Unfortunately, for technical reasons we cannot simply use the whole set-theoretic universe directly, since it is a proper class. and contains all the objects that are relevant for the proof, then construct a chain of elementary submodels of this set. Then this chain usually partitions the uncountable structure in question into pieces that inherit some important properties of the whole structure. For a gentle introduction to elementary submodels and their combinatorial applications we refer the reader to [23]. Next we recall some important facts about elementary submodels, which will allow us to flesh out the above ideas in more detail.
The transitive closure of a set $X$ is the set whose elements are: the elements of $X$, the elements of the elements of $X$ etc. For a cardinal $\varTheta$, $H(\varTheta)$ denotes the set of those sets $X$ whose transitive closure is smaller than $\Theta$. We write $\mathcal{E}\prec H(\varTheta)$ if $(\mathcal{E},\in)$ is an elementary submodel of the first order structure $(H(\Theta),\in)$, i.e. $\emptyset\neq\mathcal{E}\subseteq H(\varTheta)$ and for every formula $\varphi(v_{0},\dots,v_{n-1})$ in the language of set theory and for every $x_{0},\dots,x_{n-1}\in\mathcal{E}$: $\varphi(x_{0},\dots,x_{n-1})$ holds true in the structure $(\mathcal{E},\in)$ if and only if it holds true in $(H(\Theta),\in)$.
Fact 2.1 ([23, Claim 3.2]).
If $\varTheta$ is an infinite cardinal and $\mathcal{E}\prec H(\varTheta)$, then $\mathcal{E}$ contains all of its own finite subsets.
Fact 2.2 (implicit in [23, Corollary 2.6]).
Let $\aleph_{0}\leq\lambda<\kappa<\varTheta$ be cardinals where $\varTheta$ is regular and let $x\in H(\varTheta)$. Then there is a $\subseteq$-increasing continuous sequence $\left\langle\mathcal{E}_{\alpha}:\ \alpha<\kappa\right\rangle$ where $\mathcal{E}_{0}=\emptyset$ and $\mathcal{E}_{\alpha}\prec H(\varTheta)$ for $1\leq\alpha<\kappa$ such that $x\in\mathcal{E}_{1}$ and for every $\alpha<\kappa$: $\left|\mathcal{E}_{\alpha}\right|\cup\{\mathcal{E}_{\alpha}\}\subseteq\mathcal{E}_{\alpha+1}$ and $\left|\mathcal{E}_{\alpha}\right|=\left|\alpha\right|\cdot\lambda$.
Fact 2.3 ([23, Claim 3.7]).
Let $\varTheta$ be an uncountable cardinal and let $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$. If $X\in\mathcal{E}$ with $X\setminus\mathcal{E}\neq\emptyset$, then $\left|X\right|>\left|\mathcal{E}\right|$ and $\left|X\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$.
Fact 2.4.
Let $\varTheta$ be an uncountable cardinal, $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$ and let $\mathcal{X}\in\mathcal{E}$ be a set of countable sets. If there is a nonempty $X\in\mathcal{X}$ with $X\cap\mathcal{E}=\emptyset$, then for any maximal set $\mathcal{Y}\in\mathcal{E}$ of pairwise disjoint elements of $\mathcal{X}$ we have $\left|\mathcal{Y}\right|>\left|\mathcal{E}\right|$ and $\left|\mathcal{Y}\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$.
Proof.
If $\mathcal{Y}\subseteq\mathcal{E}$, then $\bigcup\mathcal{Y}\subseteq\mathcal{E}$ and $\mathcal{Y}\cup\{X\}$ contradicts the maximality of $\mathcal{Y}$. Therefore $\mathcal{Y}\setminus\mathcal{E}\neq\emptyset$ and we are done by applying Fact 2.3. ∎
2.2. Matroid theory
Matroids were introduced by Whitney and Nakasawa independently (see [26]) to generalize the concept of linear independence in vector spaces. In this paper we only work with the most basic concept of infinite matroids, namely finitary matroids. A more general concept of infinite matroids was discovered by Higgs [13] and rediscovered independently by Bruhn et al. [6]. For a survey about basic facts about infinite matroids we refer to [4].
An ordered pair $M=(E,\mathcal{C})$ is a finitary matroid (defined by its circuits) if $\mathcal{C}$ is a set of nonempty, pairwise $\subseteq$-incomparable finite subsets of $E$ satisfying the following circuit elimination axiom: If $C_{0},C_{1}\in\mathcal{C}$ are distinct and $e\in C_{0}\cap C_{1}$, then there is a $C\in\mathcal{C}$ with $C\subseteq(C_{0}\cup C_{1})\setminus\{e\}$. The elements of $\mathcal{C}$ are called the circuits of $M$.
Fact 2.5 (Strong circuit elimination).
If $M=(E,\mathcal{C})$ is a finitary matroid, $C_{0},C_{1}\in\mathcal{C}$ with $e\in C_{0}\setminus C_{1}$ and $f\in C_{0}\cap C_{1}$, then there is a $C\in\mathcal{C}$ with $e\in C\subseteq(C_{0}\cup C_{1})\setminus\{f\}$.
Let $M=(E,\mathcal{C})$ be a finitary matroid. Let $F$ be any set [^5] In order to simplify some later notation, we do not even assume $F\subseteq E$.. We write $M\setminus F$ for the matroid with $E(M\setminus F):=E\setminus F$ and $\mathcal{C}(M\setminus F):=\{C\in\mathcal{C}:\ C\cap F=\emptyset\}$.
Fact 2.6.
For every set $F$, the minimal elements of the set $\{C\setminus F:\ C\in\mathcal{C}\}\setminus\{\emptyset\}$ form the set of circuits of a matroid.
We write $M/F$ for the matroid on $E\setminus F$ where $\mathcal{C}(M/F)$ is the set described in Fact 2.6. Let us denote $M\setminus(E\setminus F)$ and $M/(E\setminus F)$ by $M|F$ and $M.F$ respectively. Matroids of the form $M/F\setminus G$ are called the minors of $M$. The set $\mathcal{C}^{*}$ of cocircuits of $M$ consists of the minimal nonempty subsets $D$ of $E$ for which $\left|D\cap C\right|\neq 1$ for every $C\in\mathcal{C}$.
Fact 2.7.
For every $D\in\mathcal{C}^{*}$ and for every distinct $e,f\in D$, there is a $C\in\mathcal{C}$ with $C\cap D=\{e,f\}$.
Fact 2.8.
The set $\mathcal{C}^{*}(M\setminus F)$ consists of the minimal elements of $\{D\setminus F:\ D\in\mathcal{C}^{*}\}\setminus\{\emptyset\}$ while $\mathcal{C}^{*}(M/F)=\{D\in\mathcal{C}^{*}:\ D\cap F=\emptyset\}$.
Fact 2.9.
For every $e\in E$ and every partition $X\cup Y$ of $E\setminus\left\{e\right\}$, there is either a $C\in\mathcal{C}$ with $e\in C\subseteq X\cup\left\{e\right\}$ or a $D\in\mathcal{C}^{*}$ with $e\in D\subseteq Y\cup\left\{e\right\}$.
If $G=(V,E)$ is a graph, then the edge sets of the cycles of $G$ are the circuits of a finitary matroid on $E$ which is called the cycle matroid of $G$. Its cocircuits are the bonds of $G$. A vector system with the minimal linearly dependent subsets as circuits always forms a finitary matroid. Such matroids (as well as matroids that are isomorphic to such a matroid) are called linear or representable. If the vector system in question is over $\mathbb{F}_{2}$, then it is a finitary binary matroid. The cycle matroid of a graph is a binary matroid.
Fact 2.10 ([10, Theorem 2.3]).
A finitary matroid $M$ is binary if and only if $\left|C\cap D\right|$ is even for every $C\in\mathcal{C}$ and $D\in\mathcal{C}^{*}$.
By the nature of linear independence, representable matroids are always finitary. Representability should not be confused with the more general concept of thin sum representability introduced in [1], which does not require that the matroid be finitary.
2.3. Oriented matroids
Oriented binary matroids were defined first by Minty in [18]. These are also known as signed matroids (see [21, p. 458]). The more general concept of oriented matroids was discovered independently by Bland and Las Vergnas [3] and by Folkman and Lawrence [11]. General concepts of infinite oriented matroids were introduced by Hochstättler and Kaspar [14]. Infinite binary orientable matroids are characterised (under the name of signable matroids) by Bowler and Carmesin in [5, p. 111]. Affine representable infinite oriented matroids were investigated recently by Delucchi and Knauer [8].
A signing of a set is a $\pm 1$-valued function defined on the set. A signed set is a function which is a signing of some set. A signed set is positive if it is constant $1$ and negative if it is constant 1-1. The domain of a signed set $X$ is denoted by $\underline{X}$. We define $X^{+}:=\{e\in\underline{X}:\ X(e)=1\}$ and $X^{-}:=\{e\in\underline{X}:\ X(e)=-1\}$. If $\mathcal{X}$ is a set of signed sets, then let $\underline{\mathcal{X}}:=\{\underline{X}:\ X\in\mathcal{X}\}$. A signature of a set $\mathcal{F}$ of nonempty sets is a set that consists of exactly two signings of each $F\in\mathcal{F}$ and these two are opposite. An orientation $\vec{M}$ of a finitary matroid $M=(E,\mathcal{C})$ is a pair $(E,\mathcal{O})$ where $\mathcal{O}$ is a signature of $\mathcal{C}$ that satisfies the following oriented circuit elimination axiom: For every $C_{0},C_{1}\in\mathcal{O}$ with $C_{1}\neq\pm C_{0}$ and $e\in C_{0}^{+}\cap C_{1}^{-}$, there is a $C\in\mathcal{C}$ with $C^{-}\subseteq(C_{0}^{-}\cup C_{1}^{-})\setminus\{e\}$ and $C^{+}\subseteq(C_{0}^{+}\cup C_{1}^{+})\setminus\{e\}$. Every oriented matroid gives rise to a unique signature $\mathcal{O}^{*}$ of its cocircuits satisfying
(1) $(\forall C\in\mathcal{O})(\forall D\in\mathcal{O}^{*})\left[\underline{C}\cap\underline{D}\neq\emptyset\Longrightarrow C\!\!\upharpoonright\!\!(\underline{C}\cap\underline{D})\neq\pm D\!\!\upharpoonright\!\!(\underline{C}\cap\underline{D})\right].$
Conversely, if $\mathcal{O}$ and $\mathcal{O}^{*}$ are signatures of the circuits and of the cocircuits respectively and they satisfy orthogonality axiom (1), then $\mathcal{O}$ satisfies the oriented circuit elimination axiom. The orientations of finitary matroids are called oriented finitary matroids.
Fact 2.11.
A finitary oriented matroid $\vec{M}=(E,\mathcal{O})$ is binary if and only if $\sum_{e\in\underline{C}\cap\underline{D}}C(e)D(e)=0$ for every $C\in\mathcal{O}$ and $D\in\mathcal{O}^{*}$.
The fact above is well-known in the finite case [21, Corollary 13.4.6] and remains true by compactness for finitary matroids.
We refer to the positive and negative (co)circuits of $\vec{M}$ as its directed (co)circuits.
Fact 2.12 (Farkas lemma [2, Corollary 3.4.6]).
If $\vec{M}=(E,\mathcal{O})$ is a finite oriented matroid, then for every $e\in E$ there is either a positive $C\in\mathcal{O}$ through $e$ or a positive $D\in\mathcal{O}^{*}$ through $e$ but not both.
Minors of finitary oriented matroids are defined similarly to the unoriented ones, except they inherit the signs as well.
Fact 2.13 ([14, Lemma 3.6]).
Let $\vec{M}=(E,\mathcal{O})$ be a finitary oriented matroid and let $F$ be a set. Then
$\mathcal{O}(\vec{M}\setminus F)$ $=\{C\in\mathcal{O}:\ \underline{C}\cap F=\emptyset\},$
$\mathcal{O}(\vec{M}/F)$ $=\{C\!\!\upharpoonright\!\!(E\setminus F):\ C\in\mathcal{O},\ \underline{C}\setminus F\in\mathcal{C}(M/F)\},$
$\mathcal{O}^{*}(\vec{M}\setminus F)$ $=\{D\!\!\upharpoonright\!\!(E\setminus F):\ D\in\mathcal{O}^{*},\ \underline{D}\setminus F\in\mathcal{C}^{*}(M/F)\},$
$\mathcal{O}^{*}(\vec{M}/F)$ $=\{D\in\mathcal{O}^{*}:\ \underline{D}\cap F=\emptyset\}.$
A finitary matroid is called regular if it is representable over every field. Finitary binary orientable matroids are exactly the finitary regular matroids. By replacing representability with the already mentioned thin sum representability, this connection extends even beyond finitary matroids (see [5, Subsection 5.1]).
3. Finite matching extendable matroids
Using the terminology of Komjáth et al. from [16], a matching in a matroid $M=(E,\mathcal{C})$ is a set $\mathcal{A}$ of pairwise disjoint elements of $\mathcal{C}$. A perfect matching is a matching $\mathcal{A}$ with $\bigcup\mathcal{A}=E$, i. e. a partition of the ground set of $M$ into circuits. A matroid is matchable if it admits a perfect matching. A matroid $M$ is called finite matching extendable (abbreviated as f.m.e.) if for every finite matching $\mathcal{A}$ and for every edge $e\in E\setminus\bigcup\mathcal{A}$ there is a circuit $C$ with $e\in C\subseteq E\setminus\bigcup\mathcal{A}$.
[…]
Proof.
Let $\varTheta$ be an arbitrary but fixed regular uncountable cardinal. It is sufficient to prove the theorem only for matroids in $H(\varTheta)$.
Observation 3.1.
If $\mathcal{E}\prec H(\varTheta)$, then for every finitary f.m.e. matroid $M\in\mathcal{E}$, the matroid $M|\mathcal{E}$ is also f.m.e.
Proof.
Let $\mathcal{A}$ be a finite matching in $M|\mathcal{E}$, and an let $e\in E\cap\mathcal{E}\setminus\bigcup\mathcal{A}$. Then for each $C\in\mathcal{A}$ we have $C\in\mathcal{E}$, moreover, $\mathcal{A}\in\mathcal{E}$ (see Fact 2.1). Thus $M,\mathcal{A},e\in\mathcal{E}$ and hence there is a $C\in\mathcal{C}\cap\mathcal{E}$ with $e\in C\subseteq E\setminus\bigcup\mathcal{A}$. Finally, $C\subseteq\mathcal{E}$ because $C$ is finite (see Fact 2.3), therefore $C\in\mathcal{C}(M|\mathcal{E})$. ∎
Lemma 3.2.
If $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$, then for every finitary f.m.e. matroid $M\in\mathcal{E}$, the matroid $M\setminus\mathcal{E}$ is also f.m.e.
Proof.
Let a finite matching $\mathcal{A}$ in $M\setminus\mathcal{E}$ and an edge $e\in E\setminus(\bigcup\mathcal{A}\cup\mathcal{E})$ be given. By applying the assumption that $M$ is f.m.e., we pick a $C_{0}\in\mathcal{C}$ with $e\in C_{0}\subseteq E\setminus\bigcup\mathcal{A}$ for which $F:=C_{0}\cap\mathcal{E}$ is minimal. If $F=\emptyset$, then $C_{0}\subseteq E\setminus\mathcal{E}$ and hence $C_{0}\in\mathcal{C}(M\setminus\mathcal{E})$, thus we are done. Suppose for a contradiction that $F\neq\emptyset$. Let $\mathcal{X}$ consist of the finite sets that extend $F$ to an $M$-circuit, formally $\mathcal{X}:=\{X\subseteq E\setminus F:\ F\cup X\in\mathcal{C}\}$. Since $C_{0}\setminus\mathcal{E}\in\mathcal{X}$ is clearly disjoint from $\mathcal{E}$, we conclude by applying Fact 2.4 that there is a set $\mathcal{Y}$ of pairwise disjoint elements of $\mathcal{X}$ with $\left|\mathcal{Y}\right|>\left|\mathcal{E}\right|$. Since $\left|\mathcal{E}\cup\bigcup\mathcal{A}\cup\{e\}\right|=\left|\mathcal{E}\right|$, it follows that there is an $X\in\mathcal{Y}$ that is disjoint from $\mathcal{E}\cup\bigcup\mathcal{A}\cup\{e\}$. Let $C_{1}:=F\cup X$. Since $F\neq\emptyset$ by the indirect assumption, we can pick $f\in C_{0}\cap C_{1}$. By strong circuit elimination (Fact 2.5), it follows that there is a $C\in\mathcal{C}$ with $e\in C\subseteq(C_{0}\cup C_{1})\setminus\{f\}$. But then $e\in C\subseteq E\setminus\bigcup\mathcal{A}$ with $\left|C\cap\mathcal{E}\right|\leq\left|C_{0}\cap\mathcal{E}\right|-1$ (because $C\cap\mathcal{E}\subseteq C_{0}\cap\mathcal{E}\setminus\{f\}$) contradicting the minimality of $C_{0}$. ∎
Let $M=(E,\mathcal{C})\in H(\varTheta)$ be a f.m.e. matroid. We apply transfinite induction on $\kappa:=\left|E\right|$. If $\kappa\leq\aleph_{0}$, the statement follows by a straightforward recursion. Indeed, we fix a $\kappa$-type well-order of $E$ and in step $n$ apply finite matching extendability to pick a circuit $C_{n}$ through the least element of $E\setminus\bigcup_{i<n}C_{i}$ that is disjoint from $C_{0},\dots,C_{n-1}$.
Suppose $\kappa>\aleph_{0}$. Let $\left\langle\mathcal{E}_{\alpha}:\ \alpha<\kappa\right\rangle$ be as in Fact 2.2 with $x:=M$ and $\lambda:=\aleph_{0}$. By Lemma 3.2 we know that $M\setminus\mathcal{E}_{\alpha}$ is f.m.e. for every $\alpha<\kappa$. Since $M,\mathcal{E}_{\alpha}\in\mathcal{E}_{\alpha+1}$, we have $M\setminus\mathcal{E}_{\alpha}\in\mathcal{E}_{\alpha+1}$. But then applying Observation 3.1 with $M\setminus\mathcal{E}_{\alpha}$ and $\mathcal{E}_{\alpha+1}$ for $\alpha<\kappa$, we conclude that $(M\setminus\mathcal{E}_{\alpha})|\mathcal{E}_{\alpha+1}=M|(\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha})$ is f.m.e. for every $\alpha<\kappa$. By the induction hypothesis we can pick a perfect matching $\mathcal{A}_{\alpha}$ of $M|(\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha})$ for each $\alpha<\kappa$. Clearly, $\bigcup_{\alpha<\kappa}\mathcal{A}_{\alpha}$ is a matching in $M$ and it is perfect because $E=\bigcup_{\alpha<\kappa}E\cap(\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha})$ is a partition. ∎
4. A Laviolette theorem for finitary matroids
Let $M=(E,\mathcal{C})$ be a finitary matroid and let $\lambda$ be an infinite cardinal. An $F\subseteq E$ is $\lambda$-cocircuit faithful with respect to $M$ if for every $D\in\mathcal{C}^{*}(M|F)$ with $\left|D\right|<\lambda$ we have $D\in\mathcal{C}^{*}(M)$. A $\lambda$-cocircuit faithful partition of $M$ is a partition $\{F_{i}:\ i\in I\}$ of $E$ such that
  1. (i)
    $\left|F_{i}\right|\leq\lambda$ for every $i\in I$;
  2. (ii)
    $F_{i}$ is $\lambda$-cocircuit faithful for each $i\in I$;
  3. (iii)
    For every $D\in\mathcal{C}^{*}$ with $\left|D\right|\leq\lambda$, there exists an $i\in I$ with $D\subseteq F_{i}$.
[…]
Proof.
Let $\varTheta$ be an arbitrary but fixed regular uncountable cardinal. It is sufficient to prove the theorem only for matroids in $H(\varTheta)$.
Lemma 4.1.
Let $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$. Then for every finitary matroid $M=(E,\mathcal{C})\in\mathcal{E}$ and for every $D\in\mathcal{C}^{*}$ with $D\cap\mathcal{E}\neq\emptyset$ we have: $D\subseteq\mathcal{E}$ if $\left|D\right|\leq\left|\mathcal{E}\right|$ and $\left|D\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$ if $\left|D\right|>\left|\mathcal{E}\right|$.
Proof.
We may assume that $D\setminus\mathcal{E}\neq\emptyset$ since if $D\subseteq\mathcal{E}$, then the statement holds trivially. We need to show that $\left|D\right|>\left|\mathcal{E}\right|$ and $\left|D\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$. Let $e\in D\cap\mathcal{E}$ and $f\in D\setminus\mathcal{E}$. By Fact 2.7, there is a $C\in\mathcal{C}$ with $C\cap D=\{e,f\}$. Let $F:=C\cap\mathcal{E}$ and $\mathcal{X}:=\{X\subseteq E\setminus F:\ F\cup X\in\mathcal{C}\}$. Note that each $X\in\mathcal{X}$ meets $D$ since otherwise $(F\cup X)\cap D=\{e\}$ contradicts $D\in\mathcal{C}^{*}$. Clearly, $C\setminus\mathcal{E}\in\mathcal{X}$ is disjoint from $\mathcal{E}$. Let $\mathcal{Y}\in\mathcal{E}$ be a maximal family of pairwise disjoint elements of $\mathcal{X}$. By Fact 2.4 we know that $\left|\mathcal{Y}\right|>\left|\mathcal{E}\right|$ and $\left|\mathcal{Y}\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$. Since the elements of $\mathcal{Y}$ are pairwise disjoint and each of them meets $D$, we conclude that $\left|D\right|>\left|\mathcal{E}\right|$ and $\left|D\cap\mathcal{E}\right|=\left|\mathcal{E}\right|$. ∎
Let $M=(E,\mathcal{C})\in H(\varTheta)$ be a finitary matroid. We apply transfinite induction on $\kappa:=\left|E\right|$. If $\kappa\leq\lambda$, then the singleton partition $\left\{E\right\}$ is $\lambda$-cocircuit faithful. If $\kappa>\lambda$, then let $\left\langle\mathcal{E}_{\alpha}:\ \alpha<\kappa\right\rangle$ be as in Fact 2.2 where $x=M$. By Lemma 4.1, for every $D\in\mathcal{C}^{*}$ and for the smallest ordinal $\alpha+1$ for which $D\cap\mathcal{E}_{\alpha+1}\neq\emptyset$, we have $D\subseteq\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha}$ if $\left|D\right|\leq\lambda$ and $\left|D\cap(\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha})\right|=\lambda$ if $\left|D\right|>\lambda$. We define $F_{\alpha}:=E\cap(\mathcal{E}_{\alpha+1}\setminus\mathcal{E}_{\alpha})$ for $\alpha<\lambda$. Then the partition $E=\bigcup_{\alpha<\lambda}F_{\alpha}$ satisfies (ii) and (iii), moreover $\left|F_{\alpha}\right|<\kappa$ for each $\alpha<\kappa$. Apply the induction hypothesis to $M|F_{\alpha}$ and let $F_{\alpha}=\bigcup_{i\in I_{\alpha}}F_{\alpha,i}$ be the resulting partition. We claim that $E=\bigcup_{\alpha<\lambda,i\in I_{\alpha}}F_{\alpha,i}$ is a $\lambda$-cocircuit-faithful partition of $M$. Property (i) is clear by the induction hypothesis. If $D\in\mathcal{C}^{*}(M|F_{\alpha,i})$ with $\left|D\right|<\kappa$, then by induction $D\in\mathcal{C}^{*}(M|F_{\alpha})$. But then $D\in\mathcal{C}^{*}$ because partition $E=\bigcup_{\alpha<\lambda}F_{\alpha}$ satisfies (ii). Therefore (ii) holds true. Finally, let $D\in\mathcal{C}^{*}$ with $\left|D\right|\leq\kappa$. Then there is an $\alpha$ with $D\in\mathcal{C}^{*}(M|F_{\alpha})$ because partition $E=\bigcup_{\alpha<\lambda}F_{\alpha}$ satisfies (iii). But then by induction there is an $i\in I_{\alpha}$ such that $D\in\mathcal{C}^{*}(M|F_{\alpha,i})$. Thus (iii) holds as well. ∎
5. Farkas lemma in finitary oriented matroids
First, we give a simple example due to Kaspar showing that the Farkas lemma (Fact 2.12) does not extend to finitary oriented matroids. Then we prove that it does extend for binary ones.
Claim 5.1.
There exists a finitary oriented matroid $\vec{M}=(E,\mathcal{O})$ that has neither a positive circuit nor a positive cocircuit.
Proof.
Let $\vec{M}:=(\mathbb{Z},\mathcal{O})$ where $\underline{\mathcal{O}}$ consists of the $3$-element subsets of $\mathbb{Z}$ and for $i<j<k$ the signed set $\{\left\langle i,1\right\rangle,\left\langle j,-1\right\rangle,\left\langle k,1\right\rangle\}$ and its negation are in $\mathcal{O}$. Then $\underline{\mathcal{O}^{*}}=\{\mathbb{Z}\setminus\{n\}:\ n\in\mathbb{Z}\}$ and $\mathcal{O}^{*}$ consists of the signed set $\{\left\langle m,1\right\rangle:\ m<n\}\cup\{\left\langle m,-1\right\rangle:\ n<m\}$ and its negation for each $n\in\mathbb{Z}$. It is straightforward to verify that the orthogonality axiom (1) holds, thus $\mathcal{O}$ and $\mathcal{O}^{*}$ are indeed the circuit signature and cocircuit signature of a matroid respectively. ∎
[…]
Proof.
We cannot have both by Fact 2.11. Suppose for a contradiction that there is neither a positive circuit nor a positive cocircuit through $e_{0}\in E$ in $\vec{M}$. By Fact 2.12 and by the non-existence of positive circuits through $e_{0}$, we know that for every finite $E^{\prime}\subseteq E$ containing $e_{0}$, we can find a positive cocircuit of $\vec{M}|E^{\prime}$ through $e_{0}$. Thus by compactness there exists a set $F\subseteq E$ containing $e_{0}$ such that $\sum_{e\in F\cap\underline{C}}C(e)=0$ for every $C\in\mathcal{O}$. By Zorn’s lemma we can assume that $F$ is minimal with respect to these properties. Clearly, $\left|F\cap\underline{C}\right|\neq 1$ for each $C\in\mathcal{O}$. In particular, there is no $C\in\mathcal{O}$ with $F\cap\underline{C}=\left\{e_{0}\right\}$. Thus by Fact 2.9 there is a $D\in\mathcal{O}^{*}$ with $e_{0}\in\underline{D}\subseteq F$. By negating $D$ if necessary we can assume $D(e_{0})=1$. By the indirect assumption $D$ is not a positive cocircuit, thus there is an $e_{1}\in\underline{D}$ with $D(e_{1})=-1$. Let $\mymathbb{1}_{F}$ be the characteristic function of $F$ (defined on $E$) and let $\widehat{D}$ be the extensions of $D$ to the domain $E$ with 0 values. We define $f:=\widehat{D}+\mymathbb{1}_{F}$. Then
  1. (i)
    $f\geq 0$,
  2. (ii)
    $f(e_{0})=2$,
  3. (iii)
    $\sum_{e\in\underline{C}}f(e)C(e)=0$ for every $C\in\mathcal{O}$.
For the support $S:=\{e\in E:\ f(e)\neq 0\}$ of $f$ we have $S\subseteq F\setminus\{e_{1}\}\subsetneq F$. Thus in order to get a contradiction to the minimality of $F$, we will use compactness to construct a $G$ with $e_{0}\in G\subseteq S$ such that $\sum_{e\in G\cap\underline{C}}C(e)=0$ for every $C\in\mathcal{O}$.
Let $S^{\prime}\subseteq S$ be any finite set containing $e_{0}$. If $(\vec{M}.S)|S^{\prime}$ has a positive circuit $C$ through $e_{0}$, then by Fact 2.13 there is a $C^{\prime}\in\mathcal{O}$ with $C^{\prime}\!\!\upharpoonright\!\!S=C$. But then by applying (i) and (ii) together with the positivity of $C$ we obtain
$\sum_{e\in\underline{C^{\prime}}}C^{\prime}(e)f(e)=\sum_{e\in\underline{C}}C(e)f(e)\geq f(e_{0})C(e_{0})=2,$
which contradicts (iii). Hence there is no such a circuit and therefore $(M.S)|S^{\prime}$ has a positive cocircuit $D^{\prime}$ through $e_{0}$ by Fact 2.12. Then $e_{0}\in\underline{D^{\prime}}\subseteq S$ and we claim that for every $C\in\mathcal{O}$ with $\underline{C}\cap S\subseteq S^{\prime}$ we have $\sum_{e\in\underline{D^{\prime}}\cap\underline{C}}C(e)=0$. Indeed, $D^{\prime}\in\mathcal{O}^{*}(M.S|S^{\prime})$ implies that there is a $D^{\prime\prime}\in\mathcal{O}^{*}$ with $D^{\prime\prime}\!\!\upharpoonright\!\!S^{\prime}=D^{\prime}$ and $\underline{D^{\prime\prime}}\subseteq S$ (see Fact 2.13), thus $\sum_{e\in\underline{D^{\prime}}\cap\underline{C}}C(e)=\sum_{e\in\underline{D^{\prime\prime}}\cap\underline{C}}C(e)D^{\prime\prime}(e)=0$ because $\underline{D^{\prime\prime}}\cap\underline{C}=\underline{D^{\prime}}\cap\underline{C}\subseteq S^{\prime}$. But then, since the finite set $S^{\prime}\subseteq S$ was arbitrary, it follows by compactness that the desired $G$ exists, which concludes the proof. ∎
6. Partitioning oriented matroids into directed circuits
[…]
Proof.
Let $\vec{M}=(E,\mathcal{O})$ be a finitary oriented matroid. Suppose that $\mathcal{A}$ is a set of directed circuits, say positive ones, such that $\underline{\mathcal{A}}$ is a partition of $E$. Then, for any $D\in\mathcal{O}^{*}$ and $s\in\{+,-\}$, we have $D^{s}=\dot{\bigcup}\{\underline{C}\cap D^{s}:\ C\in\mathcal{A}\}$. It follows from Fact 2.11, that we must have $\left|\underline{C}\cap D^{+}\right|=\left|\underline{C}\cap D^{-}\right|$ for every positive circuit $C$, therefore
$\left|D^{+}\right|=\sum_{C\in\mathcal{A}}\left|\underline{C}\cap D^{+}\right|=\sum_{C\in\mathcal{A}}\left|\underline{C}\cap D^{-}\right|=\left|D^{-}\right|.$
We turn to the non-trivial direction. Let a regular uncountable cardinal $\varTheta$ be fixed. We need a couple of lemmas:
Lemma 6.1.
If $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$, then for every finitary balanced oriented matroid $\vec{M}\in\mathcal{E}$, the matroid $\vec{M}\setminus\mathcal{E}$ is also balanced.
Proof.
Let $D_{0}\in\mathcal{O}^{*}(\vec{M}\setminus\mathcal{E})$ be given. By Fact 2.13 there is a $D_{1}\in\mathcal{O}^{*}(\vec{M})$ such that $D_{0}=D_{1}\!\!\upharpoonright\!\!(E\setminus\mathcal{E})$. If $\underline{D_{1}}\cap\mathcal{E}=\emptyset$, then $D_{0}=D_{1}\in\mathcal{O}^{*}(\vec{M})$ and we are done because $D_{1}$ is balanced by assumption. Suppose that $\underline{D_{1}}\cap\mathcal{E}\neq\emptyset$. Then it follows by applying Lemma 4.1 with $\underline{D_{1}}$ and $M$ that we must have $\left|\underline{D_{1}}\right|>\left|\mathcal{E}\right|$. Since $\left|D_{1}\setminus D_{0}\right|<\left|\mathcal{E}\right|$, we conclude that $\left|D_{1}^{s}\right|=\left|D_{0}^{s}\right|$ for $s\in\{+,-\}$, therefore
$\left|D_{0}^{+}\right|=\left|D_{1}^{+}\right|=\left|D_{1}^{-}\right|=\left|D_{0}^{-}\right|.$
Lemma 6.2.
If $\mathcal{E}\prec H(\varTheta)$ with $\left|\mathcal{E}\right|\subseteq\mathcal{E}$, then for every finitary binary balanced oriented matroid $\vec{M}=(E,\mathcal{O})\in\mathcal{E}$, the matroid $\vec{M}|\mathcal{E}$ is also balanced.
Proof.
Let $D\in\mathcal{O}^{*}(\vec{M}|\mathcal{E})$ be arbitrary. By symmetry we may assume that $\left|D^{+}\right|\geq\left|D^{-}\right|$. We need to show that $\left|D^{+}\right|\leq\left|D^{-}\right|$. Clearly, $D^{+}\neq\emptyset$ because $D\neq\emptyset$. Fix an $e_{0}\in D^{+}$ and let $\vec{N}:=\vec{M}/(\mathcal{E}\setminus\underline{D})\setminus(\underline{D}\setminus\{e_{0}\})$. We apply Lemma 1 with $e_{0}$ and $\vec{N}$. Suppose first that there is a positive $D_{0}\in\mathcal{O}^{*}(\vec{N})$ through $e_{0}$. Then there is a $D_{1}\in\mathcal{O}^{*}(\vec{M})$ with $D_{1}\!\!\upharpoonright\!\!((E\setminus\underline{D})\cup\{e_{0}\})=D_{0}$ and $\underline{D_{1}}\cap\mathcal{E}\subseteq\underline{D}$ (see Fact 2.13). We must have $\underline{D_{1}}\cap\mathcal{E}=\underline{D}$ because $\underline{D_{1}}\cap\mathcal{E}$ includes a cocircuit of $M|\mathcal{E}$ by Fact 2.8 but this cannot be a proper subset of another cocircuit $\underline{D}$. Then $D_{1}\!\!\upharpoonright\!\!\mathcal{E}=\pm D$ and by $D_{1}(e_{0})=D(e_{0})=1$ we conclude that $D_{1}\!\!\upharpoonright\!\!\mathcal{E}=D$. But then by the positivity of $D_{1}\setminus D$ and by the assumption that $\vec{M}$ is balanced:
$\left|D^{+}\right|\leq\left|D_{1}^{+}\right|=\left|D_{1}^{-}\right|=\left|D^{-}\right|.$
Assume now that there is a positive $C_{0}\in\mathcal{O}(\vec{N})$ through $e_{0}$. Then there is a $C_{1}\in\mathcal{O}(\vec{M})$ with $C_{1}\!\!\upharpoonright\!\!(E\setminus(\mathcal{E}\setminus\underline{D}))=C_{0}$. Let $F:=C_{1}\cap\mathcal{E}$ and let $\mathcal{X}\in\mathcal{E}$ be the set of those positive signed sets $X$ with $\underline{X}\subseteq E\setminus\underline{F}$ for which $F\cup X\in\mathcal{O}(\vec{M})$. Then $\underline{C_{1}}\setminus\mathcal{E}\in\underline{\mathcal{X}}$ is disjoint from $\mathcal{E}$ and $\underline{C_{1}}\setminus\mathcal{E}\neq\emptyset$ because $e_{0}$ cannot be an $M$-loop. It follows by Fact 2.4 that one can find $\left|\mathcal{E}\right|$ many pairwise disjoint elements in $\underline{\mathcal{X}}\cap\mathcal{E}$. To prove $\left|D^{+}\right|\leq\left|D^{-}\right|$, it suffices to show that for each $X\in\mathcal{X}\cap\mathcal{E}$, there is an $e\in\underline{D}\cap\underline{X}$ with D(e)=1D(e)=-1. Suppose that $X$ is a counterexample, and let $D^{\prime}\in\mathcal{O}^{*}(\vec{M})$ with $D^{\prime}\!\!\upharpoonright\!\!\mathcal{E}=D$. Note that $\underline{D^{\prime}}\cap(\underline{F}\cup\underline{X})=\underline{D}\cap(\underline{F}\cup\underline{X})$ because $\underline{F}\cup\underline{X}\subseteq\mathcal{E}$ and recall that $\underline{F}\cap\underline{D}=\{e_{0}\}$ by construction. Then by using the indirect assumption and the positivity of $X$:
$1=D(e_{0})F(e_{0})\leq\sum_{e\in\underline{D}\cap(\underline{X}\cup\underline{F})}D(e)(F\cup X)(e)=\sum_{e\in\underline{D^{\prime}}\cap(\underline{X}\cup\underline{F})}D^{\prime}(e)(F\cup X)(e).$
But $D^{\prime}\in\mathcal{O}^{*}(\vec{M})$ and $F\cup X\in\mathcal{O}(\vec{M})$, thus this contradicts Fact 2.11. ∎
We prove by induction on $\kappa:=\left|E\right|$ that every finitary binary balanced oriented matroid $\vec{M}=(E,\mathcal{O})\in H(\varTheta)$ admits a partition into positive circuits.
Observation 6.3.
In a finitary binary balanced oriented matroid every edge is contained in a positive circuit.
Proof.
This follows directly from Lemma 1 because being balanced ensures the non-existence of positive cocircuits. ∎
Observation 6.4.
If $\mathcal{A}$ is a finite set of pairwise disjoint positive circuits of $\vec{M}$, then $\vec{M}\setminus\bigcup\underline{\mathcal{A}}$ is balanced.
Proof.
Let $D\in\mathcal{O}^{*}(\vec{M}\setminus\bigcup\underline{\mathcal{A}})$. Pick a $D^{\prime}\in\mathcal{O}^{*}$ with $D^{\prime}\!\!\upharpoonright\!\!(E\setminus\bigcup\underline{\mathcal{A}})=D$. Since $\vec{M}$ is binary and each $C\in\mathcal{A}$ is positive, Fact 2.11 ensures that $D^{\prime}$ has the same number of positive and negative edges in $\bigcup\underline{\mathcal{A}}$. But then so does $D$ because $D^{\prime}$ is balanced by assumption. ∎
If $\kappa\leq\aleph_{0}$, then the desired partition can be obtained by a straightforward recursion via Observations 6.3 and 6.4. Indeed, we fix a $\kappa$-type well-order of $E$ and in step $n$ apply Observation 6.3 to $\vec{M}\setminus\bigcup_{i<n}\underline{C_{i}}$ (which is balanced by Observations 6.4) to find a positive circuit $C_{n}$ disjoint from $C_{0},\dots,C_{n-1}$ through the smallest element of $E\setminus\bigcup_{i<n}C_{i}$ .
Suppose $\kappa>\aleph_{0}$. Let $\left\langle\mathcal{E}_{\alpha}:\ \alpha<\kappa\right\rangle$ be as in Fact 2.2 with $x=\vec{M}$ and $\lambda=\aleph_{0}$. Then, for each $\alpha<\kappa$: $\vec{M}\setminus\mathcal{E}_{\alpha}$ is balanced by Lemma 6.1. Since $\vec{M},\mathcal{E}_{\alpha}\in\mathcal{E}_{\alpha+1}$, we have $\vec{M}\setminus\mathcal{E}_{\alpha}\in\mathcal{E}_{\alpha+1}$, thus by Lemma 6.2 we conclude that $(\vec{M}\setminus\mathcal{E}_{\alpha})|\mathcal{E}_{\alpha+1}$ is balanced. By the induction hypotheses, there is a partition $\mathcal{A}_{\alpha}$ of $(\vec{M}\setminus\mathcal{E}_{\alpha})|\mathcal{E}_{\alpha+1}$ into positive circuits. Finally, $\bigcup_{\alpha<\kappa}\mathcal{A}_{\alpha}$ is a partition of $\vec{M}$ into positive circuits. ∎
7. Open questions
A concept of infinite matroids, which is more general than finitary matroids, was introduced by Higgs and was investigated by Higgs and Oxley in the 1960s and 1970s (see [13, 20]). The investigation of this concept gained a new momentum half a century later when Bruhn et al. rediscovered it independently and provided five sets of crypcryptomorphic axiomatisations (see [6]). The axiomatisation in terms of independent sets reads as follows: $M=(E,\mathcal{I})$ is a matroid if $\mathcal{I}\subseteq\mathcal{P}(E)$ such that
  1. (1)
    $\emptyset\in\mathcal{I}$,
  2. (2)
    $\mathcal{I}$ is downward closed,
  3. (3)
    whenever $B$ is a maximal and $I$ is a non-maximal element of $\mathcal{I}$, then there is an $e\in B\setminus I$ such that $I\cup\{e\}\in\mathcal{I}$,
  4. (4)
    for every $X\subseteq E$, every $I\in\mathcal{P}(X)\cap\mathcal{I}$ can be extended to a maximal element of $\mathcal{P}(X)\cap\mathcal{I}$.
Finitary matroids are exactly those matroids in which all circuits are finite. Matroids that are not finitary are called infinitary. Relatively little is known about perfect matchings (i.e. circuit-partitions) of infinitary matroids. Next we formulate a question ispired by Theorem 2.
For an infinite cardinal $\kappa$, we call a (not necessarily finitary) matroid $\kappa$-matching extendable if for every matching (i.e. set of pairwise disjoint circuits) $\mathcal{A}$ of size less than $\kappa$ and for every edge $e\in E\setminus\bigcup\mathcal{A}$ there is a circuit $C$ through $e$ with $C\cap\bigcup\mathcal{A}=\emptyset$.
Question 7.1.
Is it true that every $\kappa$-matching extendable matroid whose circuits are smaller than $\kappa$ admits a perfect matching?
The positive answer for $\kappa=\omega$ is exactly Theorem 2.
The definition of the oriantability of infinitary matroids is somewhat problematic in general (see [14]) but works well for a certain subclass. A matroid is called tame if the intersection of any circuit with any cocircuit is finite. As in the finitary case, a tame matroid is orientable if there is a signature of its circuits and cocircuits such that they satisfy (1). We do not know whether “finitary” can be replaced by the weaker assumption “tame” in Lemma 1, i.e. the following question is open:
Question 7.2.
Is it true that in every tame binary oriented matroid $\vec{M}=(E,\mathcal{O})$, for every $e\in E$, there is either a positive circuit or a positive cocircuit through $e$?